One of the wagers in the game of roulette is to place a bet that the ball will land on a red number. (Eighteen of the numbers are black, 18 are red, and two are green.) If the ball lands on a red number, the player wins the amount of his bet. If a player bets $8, find the player's expectation. (Round your answer to two decimal places.)
Answer by Boreal(15235)

E(X)=sum of x*p(x)
=(18/38)*$8-(20/38)*8
=(144/38)-(160/38)
=-16/38 or -$0.42